Hi all I need some help with this one please. I am not sure of the result I find.

You are working on your second project as an equity research intern at a bulge investment bank. Your focus is in retail space, especially in the health and fitness sector. Currently, you are gathering information on a fast-growing chain fitness company called LuluYoga. You are interested in calculating the free cash flow of the firm.

LuluYoga offers yoga classes in several major cities in the United States. Two major revenue resources are selling workout gear and membership passes for class access.

Assume at the beginning of year 2016, LuluYoga has zero inventory.

In year 2016, LuluYoga purchased 10,000 yoga mats at a price of $10 each. The company sells 6,000 mats at a price of $15 in year 2016 and sells the remaining at a price of $20 in year 2017.

In year 2016, LuluYoga sells 1,000 membership passes for $2,000 each. 80% of the classes purchased were used in 2016 and the rest are used in 2017.The yoga master’s compensation to teach classes are $300K in year 2016 and $200K in year 2017.

LuluYoga pays corporate tax of 35%.

Question is :

What is the deferred revenue in 2016 and what is the change of NOWC in year 2016?
Answer #1

Deferred Revenue in 2016 = 0.20*1000*2000 = $ 4,00,000

Net Operating Working Capital = (Cash + Accounts Receivable + Inventories) − (Accounts Payable + Accrued Expenses)

In 2016,

Revenue : 15*6000 + 1000*2000 = $ 20,90,000

Profit before tax: $ 20,90,000-10*6000-300,000 = $ 17,30,000

PAT: 0.65* 17,30,000 = $ 11,24,500

Hence Cash =  $ 11,24,500

Inventories : 10*4000= 40,000

NOWC = 11,24,500+40,000 - 4,00,000 = $ 7,64,500